# Complete asymptotic expansions of the Humbert function $\Psi_{1}$ for two large arguments

### v3

@@ Abstract
In our recent work [SIGMA 20 (2024), 074, 13 pages], the leading behaviour of the Humbert function $\Psi_{1}[a,b;c,c^{\prime};x,y]$ when $x\to\infty$ and $y\to+\infty$ has been derived in a direct and simple manner. In this paper, we obtain the complete asymptotics of $\Psi_{1}$ in the general case $x,y\to\infty$ along a new path. Indeed, our proof is based on a sharp estimate on ${}_{2}F_{2}[a,b-n;c,d-n;z]$, which is valid uniformly for $n\in\mathbb{Z}_{\geqslant 0}$ and large $z$.
MSC 2020: primary 33C70; 41A60; secondary 30E15
Keywords: Humbert functions; Asymptotic expansions
@@ 1.  Introduction
The confluent hypergeometric function of two variables $\Psi_{1}$, introduced by Humbert [6, p. 75], is defined by
@@
By using inverse Laplace transformations of the Humbert functions $\Phi_{2}$, $\Phi_{3}$ and $\Xi_{2}$, Wald and Henkel [12] derived the leading-order behaviour of these functions when two variables become simultaneously large. But they failed to interpret the integral for $\Psi_{1}$ [12, Equation (2.4b)]
$\Psi_{1}[a,b;c,c^{\prime};x,y]=\frac{1}{\Gamma(a)}\int_{0}^{\infty}\mathrm{e}^{-u}u^{a-1}\,_{1}F_{1}\left[\begin{matrix}b\\ c\end{matrix};xu\right]{}_{0}F_{1}\left[\begin{matrix}-\\ c^{\prime}\end{matrix};yu\right]\mathrm{d}u$
@@
Assume that $a,b\in\mathbb{C},\,c,c^{\prime}\in\mathbb{C}\setminus\mathbb{Z}_{\leqslant 0}$ and $a-b,a-c,b-c\in\mathbb{C}\setminus\mathbb{Z}$. Let $w>0$ be a number such that $w>\max\{\Re(a-b),\Re(a-c)+1\}$ and that the fractional parts of $w-\Re(a-b)$ and $w-\Re(a-c)-1$ are both in the interval $(\varepsilon,1)$, where $\varepsilon>0$ is small.
@@
$\Psi_{1}={}$ $\frac{\Gamma(c)\Gamma(b-a)}{\Gamma(b)\Gamma(c-a)}A_{1}(x,y)+\frac{\Gamma(c)\Gamma(c^{\prime})\Gamma(a-b)}{\Gamma(a)\Gamma(c-b)\Gamma(b-a+c^{\prime})}(A_{2}(x,y)+A_{3}(x,y))$
$+\frac{\Gamma(c)\Gamma(c^{\prime})}{\Gamma(a)\Gamma(c-b)}A_{4}(x,y)+\mathcal{O}\left(\left|y\right|^{-\Re(b)-w}\right)+\mathcal{O}\left(\left|y\right|^{\Re(a-2b-c^{\prime})-N}\mathrm{e}^{\Re(y)}\right),$
@@
$A_{2}(x,y)={}$ $\sum_{k=0}^{M}\frac{\left(c-b\right)_{k}}{k!}\left(-y\right)^{-a-k}\sum_{j=0}^{k-1}\frac{\left(b\right)_{j}\left(a-b\right)_{k-j}\left(a-b-c^{\prime}+1\right)_{k-j}}{j!\left(c-b\right)_{k-j}}\left(\frac{y}{x-1}\right)^{b+j},$
$A_{3}(x,y)={}$ $\sum_{k=0}^{M}\frac{\left(b\right)_{k}\left(c-b\right)_{k}}{\left.k!\right.^{2}}\left(-y\right)^{-a-k}{}_{3}F_{3}\left[\begin{matrix}b+k,b-c+1,1\\ 1+k,b-a+1,b-a+c^{\prime}\end{matrix};\frac{y}{1-x}\right]\left(\frac{y}{x-1}\right)^{b+k},$
$A_{4}(x,y)={}$ $y^{a-2b-c^{\prime}}\mathrm{e}^{y}\sum_{k=0}^{N-1}a_{k}(x,y)y^{-k},$
with $M=\lfloor w+\Re(b-a)\rfloor$, $N$ being any positive integer and $a_{k}(x,y)$ given by (1.4).
@@
Moreover, the generalized hypergeometric function ${}_{p}F_{q}$ is defined by
@@
where $a_{1},\cdots,a_{p}\in\mathbb{C}$ and $b_{1},\cdots,b_{q}\in\mathbb{C}\setminus\mathbb{Z}_{\leqslant 0}$. Empty products are taken to be $1$.
@@ 2.  Auxiliary results
In this section, we prove some auxiliary results, including three lemmas about explicit bounds for the ratio of two gamma functions: (i) the first is beneficial to bound the tail series; (ii) the second generalizes the simple bound of gamma ratio [9, Equation (5.6.8)]; (iii) the third guarantees the uniformity of explicit bounds for ${}_{2}F_{2}[a,b\pm n;c,d\pm n;z]$ obtained in the ensuing theorems.
@@
Recall the integral representation for the quotient of two gamma functions [10, p. 33]: for $b>a\geqslant 0$ and $\Re(z)>0$,
@@
If $a\in\mathbb{C}$ and $b\in\mathbb{C}\setminus\mathbb{Z}$, then
@@
Assume that $a,b\in\mathbb{C},\,c\in\mathbb{C}\setminus\mathbb{Z}_{\geqslant 0}$ and $d,a-b\in\mathbb{C}\setminus\mathbb{Z}$. Let $w>0$ be a number such that $w>\max\{\Re(a),\Re(b),\Re(d)\}$ and that the fractional parts of $w-\Re(a)$ and $w-\Re(b)$ are both in the interval $(\varepsilon,1)$, where $\varepsilon>0$ is a small number. Then for any $n\in\mathbb{Z}_{\geqslant 0}$,
@@
See [7, Section 2] for details. Divide $R_{n,w}(z)$ into two parts:
@@
If $|t|\leqslant\max\{1,|\Im(d)|\}$, apply [9, Equations (5.6.6) and (5.6.7)] and (2.1) to yield
@@
Step 2. Estimate of $R_{\pm}(z)$. Due to [10, Lemma 2.1] and Euler’s reflection formula, we obtain that for $s\in L^{\pm}$,
@@
According to [10, Lemma 2.8], the contours $L^{\pm}$ can be deformed to pass through the saddle point $s=-z$ and to be bounded away from the points shown in (2.4). Hence, the integral in the last line has the order of magnitude $\mathcal{O}\big{(}|z|^{\Re(\alpha)}\mathrm{e}^{-\Re(z)}\big{)}$.
@@ 3.  Proof of the main results
For $10\leqslant n\leqslant N-$ 1, recall the exponential expansion of ${}_{2}F_{2}$ [7, Equation (5.8)]
@@
as $y\to+\infty$, where the coefficients $c_{k,n}$ are given by [11, Equation (12)]
@@
$F(x,y):=\frac{\Gamma(a)\Gamma(b)}{\Gamma(c)\Gamma(d)}\sum_{n=0}^{\infty}\frac{\left(a_{1}\right)_{n}\left(a_{2}\right)_{n}}{\left(b_{1}\right)_{n}\left(b_{2}\right)_{n}}{}_{2}F_{2}\left[\begin{matrix}a,b-n\\ c,d-n\end{matrix};y\right]\left(1-x\right)^{-n}.$
Recall the asymptotic expansion of ${}_{2}F_{2}(y)$ [7, Equation (5.8)]: for $0\leqslant n\leqslant N$,
@@
$F(x,y)=S_{w}(x,y)+T_{w}(x,y)+E_{N}(x,y)+\mathcal{O}\left(\left|y\right|^{-w}\right)+\mathcal{O}\left(\left|y\right|^{\Re(a+b-c-d)-N}\mathrm{e}^{\Re(y)}\right),$ (3.7)
@@
Dividing $T_{w}(x,y)$ into two parts:
$T_{w}(x,y)=\sum_{n=0}^{\infty}\sum_{k=0}^{n}+\sum_{n=0}^{\infty}\sum_{k=n+1}^{\lfloor w-\Re(b)\rfloor+n}=:T_{w}^{-}(x,y)+T_{w}^{+}(x,y),$
applying the transformation [8, p. 100, Equation (1)]
$\sum_{n=0}^{\infty}\sum_{k=0}^{n}A(k,n-k)=\sum_{n=0}^{\infty}\sum_{k=0}^{\infty}A(k,n)$
to $T_{w}^{-}(x,y)$ and using the series manipulation technique, we have
@@
$T_{w}^{+}(x,y)={}$ $\frac{\Gamma(b)\Gamma(a-b)}{\Gamma(c-b)\Gamma(d-b)}\sum_{k=1}^{\lfloor w-\Re(b)\rfloor}\frac{\left(b\right)_{k}\left(b-c+1\right)_{k}\left(b-d+1\right)_{k}}{\left(1\right)_{k}\left(b-a+1\right)_{k}}$
$\qquad\cdot\left(-y\right)^{-b-k}{}_{5}F_{4}\left[\begin{matrix}a_{1},a_{2},a_{3},1,b-d+k+1\\ b_{1},b_{2},b_{3},k+1\end{matrix};\frac{1}{1-x}\right]$
[…]
$T_{w}^{-}(x,y)={}$ $\frac{\Gamma(b)\Gamma(a-b)}{\Gamma(c-b)\Gamma(d-b)}\sum_{k=0}^{\infty}\frac{\left(a_{1}\right)_{k}\left(a_{2}\right)_{k}\left(a_{3}\right)_{k}\left(b-d+1\right)_{k}}{\left(b_{1}\right)_{k}\left(b_{2}\right)_{k}\left(b_{3}\right)_{k}\left(1\right)_{k}}t_{k}^{-}(x,y)\left(-y\right)^{-b-k},$
@@
$s_{k}(x,y)$ $=\sum_{j=0}^{k}\frac{\left(a_{1}\right)_{j}\left(a_{2}\right)_{j}\left(a_{3}\right)_{j}}{\left(b_{1}\right)_{j}\left(b_{2}\right)_{j}\left(b_{3}\right)_{j}}\frac{\left(a\right)_{k-j}\left(a-c+1\right)_{k-j}}{\left(1\right)_{k-j}}\left(\frac{y}{x-1}\right)^{j},$
$t_{k}^{+}(x,y)$ $=\sum_{j=0}^{k-1}\frac{\left(a_{1}\right)_{j}\left(a_{2}\right)_{j}\left(a_{3}\right)_{j}}{\left(b_{1}\right)_{j}\left(b_{2}\right)_{j}\left(b_{3}\right)_{j}}\frac{\left(b\right)_{k-j}\left(b-c+1\right)_{k-j}}{\left(b-a+1\right)_{k-j}}\left(\frac{y}{x-1}\right)^{j},$
$t_{k}^{-}(x,y)$ $={}_{5}F_{5}\left[\begin{matrix}a_{1}+k,a_{2}+k,a_{3}+k,a-b,1\\ b_{1}+k,b_{2}+k,b_{3}+k,1-b,c-b\end{matrix};\frac{y}{1-x}\right]\left(\frac{y}{x-1}\right)^{k}.$
@@ 4.  Concluding remarks
Finally, we want to make further remarks on our Theorems 1.2 and 1.3. Note that $\Psi_{1}$ has a Kummer-type transformation [3, Equation (2.54)]
@@
which is confirmed by [4, Equation (3.4)] and can be numerically verified by the integral [2, p. 115, Equation (38)]
$\Psi_{1}[a,b;c,c^{\prime};x,y]=\frac{\Gamma(c)}{\Gamma(b)\Gamma(c-b)}\int_{0}^{1}u^{b-1}\left(1-u\right)^{c-b-1}\left(1-ux\right)^{-a}{}_{1}F_{1}\left[\begin{matrix}a\\ c^{\prime}\end{matrix};\frac{y}{1-ux}\right]\mathrm{d}u,$
where $\Re(c)>\Re(b)>0$ and $(x,y)\in\mathbb{D}_{\Psi_{1}}$ (see (1.1)).

### v4

@@ Abstract
In our recent work [SIGMA 20, 074 (2024)], the leading behaviour of the Humbert function $\Psi_{1}[a,b;c,c^{\prime};x,y]$ when $x\to\infty$ and $y\to+\infty$ has been derived in a direct and simple manner. In this paper, we obtain the complete asymptotics of $\Psi_{1}$ in the general case $x,y\to\infty$ along a new path. Indeed, our proof is based on a sharp estimate on ${}_{2}F_{2}[a,b-n;c,d-n;z]$, which is valid uniformly for $n\in\mathbb{Z}_{\geqslant 0}$ and large $z$.
Mathematics Subject Classification: 33C70, 41A60, 30E15.
Keywords: Humbert functions, asymptotic expansions.
@@ 1.  Introduction
The confluent hypergeometric function of two variables $\Psi_{1}$, introduced by Humbert [7, p. 75], is defined by
@@
By using inverse Laplace transformations of the Humbert functions $\Phi_{2}$, $\Phi_{3}$ and $\Xi_{2}$, Wald and Henkel [13] derived the leading-order behaviour of these functions when two variables become simultaneously large. But they failed to interpret the integral for $\Psi_{1}$ [13, Equation (2.4b)]
$\Psi_{1}[a,b;c,c^{\prime};x,y]=\frac{1}{\Gamma(a)}\int_{0}^{\infty}\mathrm{e}^{-u}u^{a-1}\,_{1}F_{1}\left[\begin{matrix}b\\ c\end{matrix};xu\right]{}_{0}F_{1}\left[\begin{matrix}-\\ c^{\prime}\end{matrix};yu\right]\mathrm{d}u$ (1.2)
@@
Assume that $a,b\in\mathbb{C},\,c,c^{\prime}\in\mathbb{C}\setminus\mathbb{Z}_{\leqslant 0}$ and $a-b,a-c,b-c\in\mathbb{C}\setminus\mathbb{Z}$. Let $w>0$ be a number such that $w>\max\left\{\Re(a-b)+1,\Re(a-c)+2\right\}$ and that the fractional parts of $w-\Re(a-b)$ and $w-\Re(a-c)-1$ are both in the interval $(\varepsilon,1)$, where $\varepsilon>0$ is small.
@@
$\Psi_{1}={}$ $\frac{\Gamma(c)\Gamma(b-a)}{\Gamma(b)\Gamma(c-a)}A_{1}(x,y)+\frac{\Gamma(c)\Gamma(c^{\prime})\Gamma(a-b)}{\Gamma(a)\Gamma(c-b)\Gamma(b-a+c^{\prime})}A_{2}(x,y)+\frac{\Gamma(c)\Gamma(c^{\prime})}{\Gamma(a)\Gamma(c-b)}A_{3}(x,y)$
$+\mathcal{O}\left(\left|y\right|^{-\Re(b)-w}\right)+\mathcal{O}\left(\left|y\right|^{\Re(a-2b-c^{\prime})-N}\mathrm{e}^{\Re(y)}\right),$
@@
$A_{2}(x,y)={}$ $\sum_{k=0}^{M}\frac{\left(a-b\right)_{k}\left(a-b-c^{\prime}+1\right)_{k}}{k!}\,_{2}F_{2}\left[\begin{matrix}b,b-c+1-k\\ b-a+1-k,b-a+c^{\prime}-k\end{matrix};\frac{y}{1-x}\right]\left(\frac{y}{x-1}\right)^{b}\left(-y\right)^{-a-k},$
$A_{3}(x,y)={}$ $y^{a-2b-c^{\prime}}\mathrm{e}^{y}\sum_{k=0}^{N-1}a_{k}(x,y)y^{-k},$
with $M=\lfloor w+\Re(b-a)\rfloor\geqslant 1$, $N$ being any positive integer and $a_{k}(x,y)$ given by (1.5).
@@
Moreover, the generalized hypergeometric function ${}_{p}F_{q}$ is defined by [10, Equation (16.2.1)]
@@
where $a_{1},\cdots,a_{p}\in\mathbb{C}$ and $b_{1},\cdots,b_{q}\in\mathbb{C}\setminus\mathbb{Z}_{\leqslant 0}$. Empty products and sums are taken as $1$ and 00, respectively.
@@ 2.  Auxiliary results
In this section, we prove some auxiliary results, including three lemmas about explicit bounds for the ratio of two gamma functions: (i) the first is beneficial to bound the tail series; (ii) the second generalizes the simple bound of gamma ratio [10, Equation (5.6.8)]; (iii) the third guarantees the uniformity of explicit bounds for ${}_{2}F_{2}[a,b\pm n;c,d\pm n;z]$ obtained in the ensuing theorems.
@@
Recall the integral representation for the quotient of two gamma functions [11, p. 33]: for $b>a\geqslant 0$ and $\Re(z)>0$,
@@
If $a,b\in\mathbb{C}$, then
@@
Assume that $a,b\in\mathbb{C},\,c\in\mathbb{C}\setminus\mathbb{Z}_{\geqslant 0}$ and $d,a-b\in\mathbb{C}\setminus\mathbb{Z}$. Let $w>0$ be such that $w>\max\{\Re(a),\Re(b),\Re(d)\}$ and that the fractional parts of $w-\Re(a)$ and $w-\Re(b)$ are both in the interval $(\varepsilon,1)$, where $\varepsilon>0$ is a small number. Then for any $n\in\mathbb{Z}_{\geqslant 0}$,
@@
See [8, Section 2] for details. Divide $R_{n,w}(z)$ into two parts:
@@
If $|t|\leqslant\max\{1,|\Im(d)|\}$, apply [10, Equations (5.6.6) and (5.6.7)] and (2.1) to yield
@@
Step 2. Estimate of $R_{\pm}(z)$. Due to [11, Lemma 2.1] and Euler’s reflection formula, we obtain that for $s\in L^{\pm}$,
@@
According to [11, Lemma 2.8], the contours $L^{\pm}$ can be deformed to pass through the saddle point $s=-z$ and to be bounded away from the points shown in (2.4). Hence, the integral in the last line has the order of magnitude $\mathcal{O}\big{(}|z|^{\Re(\alpha)}\mathrm{e}^{-\Re(z)}\big{)}$.
@@ 3.  Proof of the main results
For $10\leqslant n\leqslant N-$ 1, recall the exponential expansion of ${}_{2}F_{2}$ [8, Equation (5.8)]
@@
as $y\to+\infty$, where the coefficients $c_{k,n}$ are given by [12, Equation (12)]
@@
$F(x,y):=\sum_{n=0}^{\infty}\frac{\left(a_{1}\right)_{n}\left(a_{2}\right)_{n}}{\left(b_{1}\right)_{n}\left(b_{2}\right)_{n}}\,_{2}F_{2}\left[\begin{matrix}a,b-n\\ c,d-n\end{matrix};y\right]\left(1-x\right)^{-n}.$
Recall the asymptotic expansion of ${}_{2}F_{2}(y)$ [8, Equation (5.8)]: for $0\leqslant n\leqslant N$,
@@
$F(x,y)=\frac{\Gamma(c)\Gamma(d)}{\Gamma(a)\Gamma(b)}\left(S_{w}(x,y)+T_{w}(x,y)+E_{N}(x,y)\right)+\mathcal{O}\left(\left|y\right|^{-w}\right)+\mathcal{O}\left(\left|y\right|^{\Re(a+b-c-d)-N}\mathrm{e}^{\Re(y)}\right),$ (3.7)
@@
Direct computation gives
@@
$s_{k}(x,y)=\sum_{j=0}^{k}\frac{\left(a_{1}\right)_{j}\left(a_{2}\right)_{j}\left(a_{3}\right)_{j}}{\left(b_{1}\right)_{j}\left(b_{2}\right)_{j}\left(b_{3}\right)_{j}}\frac{\left(a\right)_{k-j}\left(a-c+1\right)_{k-j}}{\left(1\right)_{k-j}}\left(\frac{y}{x-1}\right)^{j}.$
Write $K=\lfloor w-\Re(b)\rfloor$ and $\gamma=\frac{y}{1-x}$. Then
$T_{w}(x,y)$ $=\left(-y\right)^{-b}\sum_{n=0}^{\infty}\sum_{k=0}^{n+K}\frac{\left(a_{1}\right)_{n}\left(a_{2}\right)_{n}\left(a_{3}\right)_{n}}{\left(b_{1}\right)_{n}\left(b_{2}\right)_{n}\left(b_{3}\right)_{n}}\frac{\Gamma(b-n+k)\Gamma(a-b+n-k)}{\Gamma(d-b-k)\Gamma(c-b+n-k)}\frac{\left(-1\right)^{k}}{k!}\left(-y\right)^{n-k}\left(1-x\right)^{-n}$
$=\frac{\Gamma(b)\Gamma(a-b)}{\Gamma(c-b)\Gamma(d-b)}\left(-y\right)^{-b}\sum_{n=0}^{\infty}\sum_{k=0}^{n+K}t_{n,k}\gamma^{n}\left(-y\right)^{-k},$
where
$t_{n,k}:=\frac{\left(b\right)_{k}\left(b-c+1\right)_{k}\left(b-d+1\right)_{k}}{\left(b-a+1\right)_{k}k!}\frac{\left(a_{1}\right)_{n}\left(a_{2}\right)_{n}\left(a_{3}\right)_{n}\left(a-b-k\right)_{n}}{\left(b_{1}\right)_{n}\left(b_{2}\right)_{n}\left(b_{3}\right)_{n}\left(1-b-k\right)_{n}\left(c-b-k\right)_{n}}.$
Now split $T_{w}(x,y)$ into two parts:
$T_{w}(x,y)=\frac{\Gamma(b)\Gamma(a-b)}{\Gamma(c-b)\Gamma(d-b)}\left(-y\right)^{-b}\left(T_{w}^{-}(x,y)+T_{w}^{+}(x,y)\right),$
where
$T_{w}^{-}(x,y)$ $=\sum_{n=0}^{\infty}\sum_{k=0}^{n}t_{n,k}\gamma^{n}\left(-y\right)^{-k}=\sum_{k=0}^{\infty}\sum_{n=k}^{\infty}t_{n,k}\gamma^{n}\left(-y\right)^{-k}$
$=\sum_{k=0}^{K}\sum_{n=k}^{\infty}t_{n,k}\gamma^{n}\left(-y\right)^{-k}+\mathcal{O}\left(\left|y\right|^{-K-1}\right)$
and
$T_{w}^{-}(x,y)$ $=\sum_{n=0}^{\infty}\sum_{k=n+1}^{n+K}t_{n,k}\gamma^{n}\left(-y\right)^{-k}$
$=\sum_{n=0}^{K}\sum_{k=n+1}^{n+K}t_{n,k}\gamma^{n}\left(-y\right)^{-k}+\mathcal{O}\left(\left|y\right|^{-K-1}\right)$
$=\sum_{k=1}^{K}\sum_{n=0}^{k-1}t_{n,k}\gamma^{n}\left(-y\right)^{-k}+\mathcal{O}\left(\left|y\right|^{-K-1}\right).$
Since the inner sum in the last line is null when $k=0$, we derive
$T_{w}(x,y)$ $=\frac{\Gamma(b)\Gamma(a-b)}{\Gamma(c-b)\Gamma(d-b)}\left(-y\right)^{-b}\sum_{k=0}^{K}\sum_{n=0}^{\infty}t_{n,k}\gamma^{n}\left(-y\right)^{-k}+\mathcal{O}\left(\left|y\right|^{-\Re(b)-K-1}\right)$
$=\frac{\Gamma(b)\Gamma(a-b)}{\Gamma(c-b)\Gamma(d-b)}\sum_{k=0}^{K}t_{k}(x,y)\left(-y\right)^{-b-k}+\mathcal{O}\left(\left|y\right|^{-\Re(b)-K-1}\right),$
where
$t_{k}(x,y)=\frac{\left(b\right)_{k}\left(b-c+1\right)_{k}\left(b-d+1\right)_{k}}{\left(b-a+1\right)_{k}k!}\,_{5}F_{5}\left[\begin{matrix}a_{1},a_{2},a_{3},1,a-b-k\\ b_{1},b_{2},b_{3},1-b-k,c-b-k\end{matrix};\frac{y}{1-x}\right].$
@@ 4.  Concluding remarks
Here we would like to make further remarks on our Theorems 1.2 and 1.3. Note that $\Psi_{1}$ has a Kummer-type transformation [3, Equation (2.54)]
@@
which is confirmed by [4, Equation (3.4)] and was mentioned in [6, p. 22].
Finally, by using Mathematica 12, we provide a numerical verification of Theorems 1.2 and 1.3. The value of $\Psi_{1}=\Psi_{1}\bigl{[}1,\frac{1}{2};\frac{1}{3},\frac{1}{4};tx,ty\bigr{]}$ for large $t>0$ is evaluated by using (1.2), and the corresponding leading term is denoted by $L_{\Psi_{1}}$. Table 1 illustrates that the ratio $\frac{\Psi_{1}}{L_{\Psi_{1}}}$ approaches unity as $t\to+\infty$.
Table 1: Numerical verification of $\Psi_{1}\bigl{[}1,\frac{1}{2};\frac{1}{3},\frac{1}{4};tx,ty\bigr{]}$.
$x$ $y$ $t$ $\frac{\Psi_{1}}{L_{\Psi_{1}}}$
$-1$ $2$ $10$ $0.971796$
$-1$ $2$ $100$ $0.997355$
$-1$ $2$ $1000$ $0.999737$
$-1$ $-2$ $10$ $1.045341$
$-1$ $-2$ $100$ $1.004387$
$-1$ $-2$ $1000$ $1.000438$